OpenStudy (anonymous):

find the slope and y-intercept of the equation: 4x-3y=-6

hartnn (hartnn):

can u isolate y from that equation ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so what do u get ? y=.... ?

OpenStudy (anonymous):

4/3x+2 ???

hartnn (hartnn):

yup, now y= mx+c has slope =m and y-intercept = c

OpenStudy (anonymous):

In this equation you are trying to find y. So 4x-3y=-6 Subtract 4x from both sides then you get -3y=-4x-6 Then divide -3 from both sides then you get your answer of y=-4x-6/-3

hartnn (hartnn):

so what will be slope of y= (4/3) x + 2